Express in partial fraction x³ - x² - 4 divided by x² - 1


1
Expert's answer
2021-06-01T16:19:20-0400

"\frac{x^3-x^2-4}{x^2-1}"


"\frac{x^2(x-1)-4}{x^2-1}"


"\frac{x^2(x-1)}{(x+1)(x-1)}-\frac{4}{x^2-1}"


"x-1+\frac{x-5}{x^2-1}"


so,"\space \frac{x-5}{x^2-1}=\frac{A}{x+1}+\frac{B}{x-1}"


"=\frac{x(A+B)-A+B}{(x-1)(x+1})"


"\begin{cases}
 a &\text A+B=1 \\
 c &\text or\space -A+B=-5
\end{cases}"


Solving A and B then


"A=3\space and \space B=-2"


"so, \space \frac{x-5}{x^2-1}=\frac{3}{x+1}+\frac{-2}{x-1}"


"\frac{x^3-x^2-4}{x^2-1}=x-1+\frac{3}{x+1}-\frac{2}{x-1}"